A car travells moves a distance of 200m.it cover first half distance at speed 40km/h and second half at speed of v km/h.calculate average speed of car


1
Expert's answer
2021-07-12T08:40:33-0400

First time"t_1=\frac{100}{40}"

Second time"t_2=\frac{100}{v}"

Total time

"t=t_1+t_2"

"t=\frac{100}{40}+\frac{100}{v}"


Average speed="\frac{total distance}{total time}"

"48=\frac{200}{\frac{100}{40}+\frac{100}{v}}"

"\frac{1}{4}+\frac{10}{v}=\frac{20}{48}"

"\frac{1}{4}+\frac{10}{v}=\frac{20}{48}"

"v=60km\/h"